OpenStudy (anonymous):

Point A is located at (0, 4), and point B is located at (−2, −3). Find the x value for the point that is the distance from point A to point B. −1 −0.75 −0.5 −0.25

OpenStudy (anonymous):

@jim_thompson5910 @shamim @TSwizzle @mathslover

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

is the questions supposed to say halfway between A and B?

OpenStudy (anonymous):

Oops sorry! it's 1/4

OpenStudy (anonymous):

 so the change in x position between A and B is -2 - 0 = -2 If you want a quarter of that distance, its -2/4 = -1/2 = -.5 Now to find the x position, add that to your starting point (in this case, A's x value is at 0) 0 + -.5 = -.5 C is your answer

OpenStudy (anonymous):

Oh, wow! Thank you so much! :) I understand this now!!